OpenStudy (anonymous):

can someone help out with algebra 1 work?????

OpenStudy (mathmale):

sashajj: would you please post your questions first and ask the "can someone help" question later on?

OpenStudy (anonymous):

choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (anonymous):

y = one thirdx - 8 y = -3x + 12 y = 3x - 24 y = -one thirdx - 4

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

anyone????

OpenStudy (anonymous):

did you graph all of the answer options??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay which one did you think it was?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

i think you are correct i didn't read the question right sorry

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

one second @whpalmer4 is this correct??

OpenStudy (anonymous):

im pretty sure that its the second one

OpenStudy (anonymous):

if i get this wrong im probably screwed because have 5 questions and i need all right and im not sure with the other once

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1 sec

OpenStudy (whpalmer4):

slope of the line in the figure is 3 slope of the perpendicular line will be -1/3 line with slope -1/3 through point (6,-6) will be \[y-(-6) = -\frac{1}{3}(x-6)\] \[y + 6 = -\frac{1}{3}x+2\]\[y=-\frac{1}{3}x-4\]

OpenStudy (anonymous):

Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations. Equation #1 = y - 4 = 1(x - 0) Equation #2 = y - 3 = 1(x - 1)

OpenStudy (anonymous):

1.Only equation #1 represents the line that passes through the two given points. 2.Only equation #2 represents the line that passes through the two given points. 3.Both equation #1 and equation #2 represent the line that passes through the two given points. 4.Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (whpalmer4):

For that first question, the 4th answer choice is the correct one, not the 2nd. To consider the second question, the form they are using is \[y - y_1 = m(x-x_1)\]Giving the equation of a line with slope \(m\) passing through point \((x_1,y_1)\)

OpenStudy (whpalmer4):

You should calculate the slope of the line passing through (0,4) and (-1,3) as a first step. Is it what the two equations assume it is?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Good. If it isn't, then we'd obviously choose answer #4.

OpenStudy (whpalmer4):

So are the two equations provided the equations you would get if you plugged m = 1 and the two points given into my point-slope formula?

OpenStudy (whpalmer4):

I encourage you to examine them very carefully before answering :-)

OpenStudy (anonymous):

yes they are the equations u would get if plugged m = 1

OpenStudy (whpalmer4):

I'm afraid not... \((0,4)=(x_1,y_1)\) \[y-y_1=m(x-x_1)\]\[y-4=1(x-0)\] \((-1,3)=(x_1,y_1)\) \[y-y_1 = m(x-x_1)\]\[y-3 = 1(x-(-1))\] Still thinking that is the same equation they gave you?

OpenStudy (anonymous):

im so confused

OpenStudy (mathmale):

@sashajj: Are you still focused on the first of these 5 problems? In other words, given the graph, are you supposed to find the equation of a line perpendicular to the given (graphed) line and passing through the point (6,-6)?

OpenStudy (anonymous):

ya i have been doing this for the past 3 hours and im not even sure if i did it right im confused with the last question i wrote here and i need 20 shots

OpenStudy (mathmale):

Let's give it another try. First I need to be sure that you know two key facts. 1) if two lines are perpendicular, their slopes are negative reciprocals of one another. Example: if the slope of that line in the illustration is 3, then the slope of a line perpendicular to it is -1/3, which is the "negative reciprocal" of 3. OK so far?

OpenStudy (mathmale):

2. Given the slope of a straight line and one point on the line, the equation of the line is \[y-y _{0}=m(x-x _{0)}\]

OpenStudy (mathmale):

@sashajj: working on a different problem right now? If so, come back here when you can; we're not that far from finishing this problem!